OpenStudy (anonymous):

A library is having a book-sale to raise money. $4 for each hardcover. $2 for paperback books A person spends $26 for 8 books. How many hardcover books did the person bought?

OpenStudy (ajprincess):

Let the number of hardcover books the person bought be x nd number of paperback books be y. Then v hav x+y=8 4x+2y=26 Solvng the simultaneous equations we can find the number of hardcover books

OpenStudy (anonymous):

let x is number of hardcover books and y is number of paperback book then x+y = 8--->A 4x+2y = 26-->B from A we have x = 8-y put this value of x in B 4(8-y)+2y = 28 can you find y now???